Subject: Raith Quirks Alert
From: Charis Quay Huei Li <>
Date: Sun, 6 Jul 2003 03:06:53 -0400 (EDT)

1) The CCD camera is no longer working. This problem first reported in the
logbook on the 4th of July by cshen. Apprently restarting the comptuer
sometimes fixes this but I decided not to try it since I could find my
sample anyway.

2) The stage has noticeable drift. About 0.1mm/hour by my estimate. Both
cshen and I have found that turning the laser stage off and on again
sometimes fixes this problem.


P.S. Perhaps there needs to be space in the logbook for reporting this
sort of thing.